OpenStudy (anonymous):

Use the sum or different to find he exact value of each trigonometric function a) tan(π)/(12)

OpenStudy (across):

Do you know what the period of the sin function is?

OpenStudy (anonymous):

note that \[\frac{\pi}{12}=\frac{\pi}{3}-\frac{\pi}{4}\] and use the subtraction angle formula for tangent

OpenStudy (across):

Remember,\[\tan(x)=\frac{\sin(x)}{\cos(x)}.\]

OpenStudy (across):

Oh, the asker meant\[\tan(\frac{\pi}{12}).\]

OpenStudy (anonymous):

\[\tan(\alpha - \beta)=\frac{\tan(\alpha)-\tan(\beta)}{1-\tan(\alpha)\tan(\beta)}\]

OpenStudy (anonymous):

so you get \[\tan(\frac{\pi}{3}- \frac{\pi}{4})=\frac{\tan(\frac{\pi}{3})-\tan(\frac{\pi}{4})}{1-\tan(\frac{\pi}{3})\tan(\frac{\pi}{4})}\]